hey,the question is_
if a sequence follows the same patterns as a, ar, ar^2... and t4=8 and t7=1
find a, r, and t10

Answer by stanbon(75887)

if a sequence follows the same patterns as a, ar, ar^2... and t4=8 and t7=1
find a, r, and t10
-------------------------------
t(4) = 8 = ar^3
t(7) = 1 = ar^6
====================
Divide t(7) by t(4) to get:
[ar^6]/[ar^3] = 1/8
r^3 = 1/8
r = 1/2
---------------
Substitute into ar^3 = 8 to solve for "a":
a(1/2)^3 = 8
a(1/8) = 8
a = 64
===============
So, t(10) = ar^9 = 64(1/2)^9 = 1/8
========================================
Cheers,
Stan H.